# Applying statistical learning theory to deep learning

### v1

@@ 5.2 Reminder on the kernel regime
In what follows, we will sometimes write $\phi_{0}(\mathbf{x})=\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x})$. In certain regimes, this linear approximation is always valid across training and the model behaves as an affine model $f(\mathbf{w},\mathbf{x})=f_{0}(\mathbf{x})+\langle\mathbf{w},\phi_{0}(\mathbf{x})\rangle$ with feature map $\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x})$ corresponding to the tangent kernel $K_{0}(\mathbf{x},\mathbf{x}^{\prime})=\langle\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x}),\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x}^{\prime})\rangle$. GD then learns the corresponding minimum RKHS distance to the initialization $F(\mathbf{w}_{0})$ solution, i.e. $\operatorname*{arg\,min}_{h}\lVert h-F(\mathbf{w}_{0})\rVert_{K_{0}}\thickspace\mbox{s.t.}\thickspace h(X)=\mathbf{y}$. We can avoid the $F(\mathbf{w}_{0})$ by choosing a familiy of functions verifying $F(\mathbf{w}_{0})=0$ (unbiased initialization from [8]). Then are we just studying an uninformative regime or can we really replace neural nets with linear models ? In what case do we get this kernel regime ?
Initially, was linked to the width, taking width to infinity leads to a Gaussian process which is a kernel regime. See e.g. [13, 9]. Closer to what we want to do is [8]. Regardless of the width, can always reach the kernel regime when the scale of the initialization goes to infinity. In what follows, we will mainly consider gradient flow, i.e.
@@ 5.3.1 Analytical study of GD in parameter space
Applying the chain rule gives, denoting $\mathbf{X}\in\mathbb{R}^{n\times d}$ the design matrix
@@
$\dot{\mathbf{w}}_{-}=-\frac{d\boldsymbol{\beta}}{d\mathbf{w}^{-}}\frac{dL(\boldsymbol{\beta})}{d\boldsymbol{\beta}}=2\mathbf{X}^{\top}\mathbf{r}(t)\odot\mathbf{w}_{-}(t)$ (67)
@@
Consider that the residuals are known, then these are differential equations that can be solved. Integrating yields
@@
where sinh is the hyperbolic sine. We are interested in the regime where $n<<d$ i.e. number of samples much lower than dimensionality $d$, so they are many solutions to the problem $\mathbf{X}\boldsymbol{\beta}=\mathbf{y}$. However, Eq.(71)-(72) shows that we have reduced the set of solutions to a lower dimensional manifold of dimension $n$. This is similar to the study of GD on linear regression, i.e. the predictor is always spanned by the data. Here we observe the same thing on a non-linear model. This manifold is defined by
@@
where we have $n$ additional constraints defined by the equation $\mathbf{X}\boldsymbol{\beta}=\mathbf{y}$ leading to a unique solution. Note that we are not studying the convergence of gradient flow, we assume it converges. Let’s write an equivalent optimization problem giving the same set of solutions. We want to find a function $Q_{\alpha}(\boldsymbol{\beta})$ that is implictly minimized by the GF dynamics such that
@@
where the inverse hyperbolic sine is applied elementwise. Matching this with the optimality condition then gives
@@
For small values of $\alpha$ however, $q$ becomes close to a $\ell_{1}$ norm and we obtain an effective $\ell_{1}$, sparsity inducing implicit regularization which does not correspond to a kernel regime, but a rich regime.
@@
We thus have a transition from kernel inductive bias to a sparsity inducing inductive bias. The rich learning regime corresponds to the $\ell_{1}$ regime, which corresponds to learning : learning feature corresponds to selecting a small number of features among an infinite amount of features. A more precise characterization of this transition can be found in Theorem 2 from [33], which we reproduce here.
@@
$y_{i}=\langle\boldsymbol{\beta}^{*},\mathbf{x}{i}\rangle+\gamma\quad\mbox{where}\quad\gamma\sim\mathcal{N}(0,0.01)$ (87)
@@
Getting to the $\ell_{1}$ regime is actually quite difficult. See this from Theorem 2 : $\alpha$ has to be exponentially small. What’s the sample complexity as a function of $\alpha$, or conversely, how small does $\alpha$ has to be to reach good performance, let’s say $L(\boldsymbol{\beta}_{\alpha}(\infty))\leqslant 0.025$
@@ 5.3.2 Studying the dynamics in function space
Whats do the dynamics look like in function space ? Recall that the function space is parametrized by $\boldsymbol{\beta}$, thus we want to write the solution directly on $\boldsymbol{\beta}$ instead of $\mathbf{w}$, and $\boldsymbol{\beta}=F(\mathbf{w})$.
@@
$=\nabla F(\mathbf{w}(t))^{\top}\left(-\nabla_{\mathbf{w}}L(\mathbf{w}(t))\right)$ (89)
@@
$\dot{\boldsymbol{\beta}}=-\nabla F(\mathbf{w}(t))^{\top}\nabla F(\mathbf{w}(t))\nabla L(\boldsymbol{\beta})$ (92)
which corresponds to what the previously discussed Riemanian gradient flow, with metric tensor $\rho=\left(\nabla F(\mathbf{w}(t))^{\top}\nabla F(\mathbf{w}(t))\right)^{-1}$, i.e.
@@
$\rho(\mathbf{w}(t))$ $=\mbox{diag}\left(\mathbf{w}_{+}^{2}+\mathbf{w}_{-}^{2}\right)^{-1}$ (97)
@@
$w_{-}^{2}=\frac{-\beta+\sqrt{\beta^{2}+4\alpha^{4}}}{2}$ (103)
@@
$w_{+}^{2}=\frac{\beta+\sqrt{\beta^{2}+4\alpha^{4}}}{2}$ (104)
Which leads to the following expression for the metric tensor $\rho$ and the corresponding dynamics
@@
We can recover the previously discussed phenomenology from this equation. For large $\alpha$, the $\beta^{2}$ term is negligible and we recover standard gradient flow dynamics with $\ell_{2}$ geometry in the function space. If $\alpha=0$, the scaling in front of the gradient is proportional to the absolute value of $\beta$. Thus higher absolute value coefficients will decay more, which will promote sparsity. Now, does this metric tensor correspond to a Hessian map defining a mirror descent ? To check this we need to solve $\rho=\nabla^{2}\Psi$ where $\Psi$ is the potential defining the Bregman distance used for mirror descent. In general, a metric tensor is not a Hessian map, but here it is the case, mostly thanks to the diagonal structure. We can then simply integrate each element on the diagonal twice. Performing this double integral yields the following potential
@@
Up to a constant, this is the same potential as the one implicitly being minimized by the gradient descent.
@@ 5.3.3 Comparing explicit and implicit regularization
Is this equivalent to explicity regularization using the $\ell_{2}$ norm ?
@@ 5.4 The effect of width
thus the entire dynamics is described by $\mathbf{M}_{\mathbf{W}\mathbf{W}^{\top}}$. In the asymmetric case $\mathbf{M}_{{\mathbf{U}},\boldsymbol{V}}$ this is not true. We may then consider the following ”lifted” problem defined by
@@
and the corresponding ”lifted” datapoints $\bar{\boldsymbol{X}}_{n}=\frac{1}{2}\begin{bmatrix}0&\boldsymbol{X}_{n}\\ \boldsymbol{X}_{n}^{\top}&0\end{bmatrix}$, where we consider that the datapoints are matrices in $\mathbb{R}^{d\times d}$, not necessarily diagonal. The implemented function is
@@
the output of which is the same as the original model but now $\bar{{\mathbf{M}}}_{{\mathbf{U}},\boldsymbol{V}}$ is the relevant matrix to study the problem. Assume that ${\mathbf{U}}(0),\boldsymbol{V}(0)$ are initialized with $\mathcal{N}(0,\sigma^{2}=\frac{\alpha^{2}}{\sqrt{k}})$. This way $\mbox{Var}\left[\mbox{diag}({\mathbf{U}}\boldsymbol{V}^{\top})[i]\right]=\alpha^{2}$. In the case where the measurements commute, the following theorem is proven in [33] (we note that the definition of scaling parameters are different from the paper here, but ultimately the statements are equivalent)
@@
[…]
@@
  • $\sqrt{k}\alpha^{2}\to 0$ leads to the kernel regime, even if $\lVert\boldsymbol{\beta}(0)\rVert\simeq\alpha^{2}\to 0$
@@ 5.5 Deep diagonal networks
$\boldsymbol{\beta}(t)={\mathbf{w}}_{+}(t)^{D}-{\mathbf{w}}_{-}(t)^{D}\quad\mbox{and}f_{D}({\mathbf{w}},\boldsymbol{x})=\langle{\mathbf{w}}_{+}(t)^{D}-{\mathbf{w}}_{-}(t)^{D},\boldsymbol{x}\rangle$ (119)
@@
where $h_{D}=\alpha^{D}\left(1+\alpha^{D-2}D(D-2)z\right)^{-\frac{D}{D-2}}-\alpha^{D}\left(1+\alpha^{D-2}D(D-2)z\right)^{-\frac{D}{D-2}}$ (we note that the function with different coefficients in the paper, but ultimately the statements are equivalent). Recall that we are searching for an equivalent problem of the form
${\boldsymbol{\beta}}^{*}\in\inf_{{\boldsymbol{\beta}}}Q({\boldsymbol{\beta}})\thickspace\mbox{s.t.}\thickspace\boldsymbol{X}{\boldsymbol{\beta}}={\mathbf{y}}$ (127)
@@
$\mathcal{L}({\boldsymbol{\beta}},\boldsymbol{\nu})=Q({\boldsymbol{\beta}})+{\boldsymbol{\nu}}^{\top}\left(\boldsymbol{X}{\boldsymbol{\beta}}-{\mathbf{y}}\right)$ (128)
@@
$\boldsymbol{X}{\boldsymbol{\beta}}={\mathbf{y}}\quad\mbox{and}\quad\nabla Q({\boldsymbol{\beta}}^{*})=\boldsymbol{X}^{\top}\boldsymbol{\nu}$ (129)
. We can then match $\mathbf{s}$ with ${\boldsymbol{\nu}}$ and identify the potential $Q$ by matching its gradient with the inverse of $h_{D}$. Then, define $q_{D}=\int h^{-1}_{D}$ and $Q_{D}({\boldsymbol{\beta}})=\sum_{i}q_{D}\left(\frac{\beta[i]}{\alpha^{D}}\right)$. It is proven in [33] that
@@
so the domain of $h_{D}$ is the interval $[-1,1]$ upon which it is monotonically incresing, ensuring the existence of the inverse mapping $h^{-1}_{D}$. Then, for all depth $D\geqslant 2$, this equivalent cost induces a rich implicit bias for $\alpha\to 0$, i.e.
$\lim_{\alpha\to 0}{\boldsymbol{\beta}}^{\infty}_{\alpha,D}={\boldsymbol{\beta}}_{\ell_{1}}^{*}\\ \lim_{\alpha\to\infty}{\boldsymbol{\beta}}^{\infty}_{\alpha,D}={\boldsymbol{\beta}}_{\ell_{2}}^{*}$ (131)
Although the same behaviour is observed as for the $D=2$ case, there are actually two main differences. The first one is that, for $D>2$, explicit regularization does not lead to a sparse bias. Indeed, for
@@
$R_{\alpha}({\boldsymbol{\beta}})\xrightarrow[]{\alpha\to 0}\lVert{\boldsymbol{\beta}}\rVert_{2/D}$ (133)

### v2

@@ 5.2 Reminder on the kernel regime
In what follows, we will sometimes write $\phi_{0}(\mathbf{x})=\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x})$. In certain regimes, this linear approximation is always valid across training and the model behaves as an affine model $f(\mathbf{w},\mathbf{x})=f_{0}(\mathbf{x})+\langle\mathbf{w},\phi_{0}(\mathbf{x})\rangle$ with feature map $\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x})$ corresponding to the tangent kernel $K_{0}(\mathbf{x},\mathbf{x}^{\prime})=\langle\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x}),\nabla_{\mathbf{w}}f(\mathbf{w}_{0},\mathbf{x}^{\prime})\rangle$. GD then learns the corresponding minimum RKHS distance to the initialization $F(\mathbf{w}_{0})$ solution, i.e. $\operatorname*{arg\,min}_{h}\lVert h-F(\mathbf{w}_{0})\rVert_{K_{0}}\thickspace\mbox{s.t.}\thickspace h(X)=\mathbf{y}$. Note that we can avoid the $F(\mathbf{w}_{0})$ by choosing a familiy of functions verifying $F(\mathbf{w}_{0})=0$ (see e.g. the unbiased initialization from [16]). Then are we just studying an uninformative regime or can we really replace neural nets with linear models? In what case does this kernel regime appear?
Initially, the appearance of the kernel regime was shown to be linked to the width of the network : taking the width to infinity under an appropriate scaling of the weights allows to linearize the network to obtain an asymptotically equivalent (in law) Gaussian process, leading to a kernel method and thus to the kernel regime. For the corresponding result, see e.g. [26, 18]. Closer to what we want to do is [16]. Regardless of the width, can always reach the kernel regime when the scale of the initialization goes to infinity. In what follows, we will mainly consider gradient flow, i.e.
@@ 5.3.1 Analytical study of GD in parameter space
Applying the chain rule gives, denoting $\mathbf{X}\in\mathbb{R}^{n\times d}$ the design matrix,
@@
$\dot{\mathbf{w}}_{-}=-\frac{d\boldsymbol{\beta}}{d\mathbf{w}^{-}}\frac{dL(\boldsymbol{\beta})}{d\boldsymbol{\beta}}=2\mathbf{X}^{\top}\mathbf{r}(t)\odot\mathbf{w}_{-}(t),$ (66)
@@
Assume that the residuals are known, then these are differential equations that can be solved. Integrating yields
@@
where sinh is the hyperbolic sine. We are interested in the regime where $n<<d$ i.e. where the number of samples is much lower than the dimensionality $d$, so they are many solutions to the problem $\mathbf{X}\boldsymbol{\beta}=\mathbf{y}$. However, Eq.(69)-(70) shows that we have reduced the set of solutions to a lower dimensional manifold of dimension $n$. This is similar to the study of GD on linear regression, i.e. the predictor is always spanned by the data. Here we observe the same thing on a non-linear model. This manifold is defined by
@@
where we have $n$ additional constraints defined by the equation $\mathbf{X}\boldsymbol{\beta}=\mathbf{y}$ leading to a unique solution. Recall that we are not studying the convergence of gradient flow per se, rather we assume it converges and we study the corresponding fixed point to characterize its implicit bias. Let’s write an equivalent optimization problem giving the same set of solutions. We want to find a function $Q_{\alpha}(\boldsymbol{\beta})$ that is implictly minimized by the GF dynamics such that
@@
where the inverse hyperbolic sine is applied elementwise. Matching this with the optimality condition Eq.(77) then gives
@@
For small values of $\alpha$ however, $q$ becomes close to a $\ell_{1}$ norm and we obtain an effective $\ell_{1}$, sparsity inducing implicit regularization which does not correspond to a kernel regime, but a rich regime :
@@
We thus have a transition from a kernel inductive bias to a sparsity inducing inductive bias. The rich learning regime corresponds to the $\ell_{1}$ regime, which is one of the main benefits we are looking for in a machine learning model : learning features corresponds to selecting a small number of important features among an infinite amount of features (in the infinite size limit for neural networks). A more precise characterization of this transition can be found in Theorem 2 from [53], which we reproduce here.
@@
$y_{i}=\langle\boldsymbol{\beta}^{*},\mathbf{x}_{i}\rangle+\gamma\quad\mbox{where}\quad\gamma\sim\mathcal{N}(0,0.01)$ (85)
@@
Getting to the $\ell_{1}$ regime is actually quite difficult. Indeed, looking at the thresholds given by Theorem 5.1 : $\alpha$ has to be exponentially small. What’s the sample complexity as a function of $\alpha$, or conversely, how small does $\alpha$ has to be to reach good performance, let’s say $L(\boldsymbol{\beta}_{\alpha}(\infty))\leqslant 0.025$
@@ 5.3.2 Studying the dynamics in function space
Whats do the dynamics look like in function space ? Recall that the function space is parametrized by $\boldsymbol{\beta}$, thus we want to write the solution directly as a function of $\boldsymbol{\beta}$ instead of $\mathbf{w}$, where $\boldsymbol{\beta}=F(\mathbf{w})$. Writing the dynamics at the level of $\beta$, we reach
@@
$=\nabla F(\mathbf{w}(t))^{\top}\left(-\nabla_{\mathbf{w}}L(\mathbf{w}(t))\right),$ (87)
@@
$\dot{\boldsymbol{\beta}}=-\nabla F(\mathbf{w}(t))^{\top}\nabla F(\mathbf{w}(t))\nabla L(\boldsymbol{\beta}).$ (90)
This corresponds to the previously discussed Riemanian gradient flow (or natural gradient descent), with a metric tensor determined by $\rho=\left(\nabla F(\mathbf{w}(t))^{\top}\nabla F(\mathbf{w}(t))\right)^{-1}$, i.e.
@@
$\rho(\mathbf{w}(t))$ $=\mbox{diag}\left(\mathbf{w}_{+}^{2}+\mathbf{w}_{-}^{2}\right)^{-1}.$ (93)
@@
$w_{-}^{2}=\frac{-\beta+\sqrt{\beta^{2}+4\alpha^{4}}}{2}.$ (99)
@@
$w_{+}^{2}=\frac{\beta+\sqrt{\beta^{2}+4\alpha^{4}}}{2}.$ (100)
This leads to the following expression for the metric tensor $\rho$ and the corresponding dynamics
@@
We can recover the previously discussed phenomenology from this equation. For large $\alpha$, the $\beta^{2}$ term is negligible and we recover standard gradient flow dynamics with $\ell_{2}$ geometry in the function space. If $\alpha=0$, the scaling in front of the gradient is proportional to the absolute value of $\beta$. Thus higher absolute value coefficients will decay faster, which will promote sparsity. Now, does this metric tensor correspond to a Hessian map defining a mirror descent? To check this we need to solve $\rho=\nabla^{2}\Psi$ where $\Psi$ is the potential defining the Bregman distance used for mirror descent. In general, a metric tensor is not a Hessian map, but here it is the case, mostly thanks to the diagonal structure. We can then simply integrate each element on the diagonal twice. Performing this double integral yields the following potential
@@
Up to a constant, this is the same potential as the one implicitly being minimized by the gradient descent, as established at Eq.82.
@@ 5.3.3 Comparing explicit and implicit regularization
Is what we have established in the previous section equivalent to explicit regularization using the $\ell_{2}$ norm in parameter space?
@@ 5.4 The effect of width
thus the entire dynamics is described by $\mathbf{M}_{\mathbf{W}\mathbf{W}^{\top}}$. In the asymmetric case $\mathbf{M}_{{\mathbf{U}},\boldsymbol{V}}$ this is not true. We may then consider the following lifted problem defined by
@@
and the corresponding lifted datapoints $\bar{\boldsymbol{X}}_{n}=\frac{1}{2}\begin{bmatrix}0&\boldsymbol{X}_{n}\\ \boldsymbol{X}_{n}^{\top}&0\end{bmatrix}$, where we consider that the datapoints are matrices in $\mathbb{R}^{d\times d}$, not necessarily diagonal. The implemented function is
@@
the output of which is the same as the original model but now $\bar{{\mathbf{M}}}_{{\mathbf{U}},\boldsymbol{V}}$ is the relevant matrix to study the problem. Assume that ${\mathbf{U}}(0),\boldsymbol{V}(0)$ are initialized with $\mathcal{N}(0,\sigma^{2}=\frac{\alpha^{2}}{\sqrt{k}})$. This way $\mbox{Var}\left[\mbox{diag}({\mathbf{U}}\boldsymbol{V}^{\top})[i]\right]=\alpha^{2}$. In the case where the measurements commute, the following theorem is proven in [53] (we note that the definition of the scaling parameters are different in [53], but ultimately the statements are equivalent)
@@
${\mathbf{M}}_{{\mathbf{U}},\boldsymbol{V}}^{*}=\operatorname*{arg\,min}_{\mathbf{M}}Q_{\mu}(\mbox{spectrum}({\mathbf{M}}))\thickspace\mbox{s.t.}\thickspace\mathbf{L}({\mathbf{M}})=0$ (114)
@@
  • the scaling $\sqrt{k}\alpha^{2}\to 0$ leads to the kernel regime, even if $\lVert\boldsymbol{\beta}(0)\rVert\simeq\alpha^{2}\to 0$
@@ 5.5 Deep diagonal networks
$\boldsymbol{\beta}(t)={\mathbf{w}}_{+}(t)^{D}-{\mathbf{w}}_{-}(t)^{D}\quad\mbox{and}\quad f_{D}({\mathbf{w}},\boldsymbol{x})=\langle{\mathbf{w}}_{+}(t)^{D}-{\mathbf{w}}_{-}(t)^{D},\boldsymbol{x}\rangle$ (115)
@@
where $h_{D}=\alpha^{D}\left(1+\alpha^{D-2}D(D-2)z\right)^{-\frac{D}{D-2}}-\alpha^{D}\left(1+\alpha^{D-2}D(D-2)z\right)^{-\frac{D}{D-2}}$ (we note that the chosen scaling is slightly different in [53], but ultimately the statements are equivalent). Recall that we are searching for an equivalent problem of the form
${\boldsymbol{\beta}}^{*}\in\inf_{{\boldsymbol{\beta}}}Q({\boldsymbol{\beta}})\thickspace\mbox{s.t.}\thickspace\boldsymbol{X}{\boldsymbol{\beta}}={\mathbf{y}},$ (123)
@@
$\mathcal{L}({\boldsymbol{\beta}},\boldsymbol{\nu})=Q({\boldsymbol{\beta}})+{\boldsymbol{\nu}}^{\top}\left(\boldsymbol{X}{\boldsymbol{\beta}}-{\mathbf{y}}\right),$ (124)
@@
$\boldsymbol{X}{\boldsymbol{\beta}}={\mathbf{y}}\quad\mbox{and}\quad\nabla Q({\boldsymbol{\beta}}^{*})=\boldsymbol{X}^{\top}\boldsymbol{\nu}.$ (125)
We can then match $\mathbf{s}$ with ${\boldsymbol{\nu}}$ and identify the potential $Q$ by matching its gradient with the inverse of $h_{D}$. To do so, define $q_{D}=\int h^{-1}_{D}$ and $Q_{D}({\boldsymbol{\beta}})=\sum_{i}q_{D}\left(\frac{\beta[i]}{\alpha^{D}}\right)$. It is proven in [53] that
@@
so the domain of $h_{D}$ is the interval $[-1,1]$ upon which it is monotonically increasing, ensuring the existence of the inverse mapping $h^{-1}_{D}$. Then, for all depth $D\geqslant 2$, this equivalent cost induces a rich implicit bias for $\alpha\to 0$, i.e.
$\lim_{\alpha\to 0}{\boldsymbol{\beta}}^{\infty}_{\alpha,D}={\boldsymbol{\beta}}_{\ell_{1}}^{*}$ (127)
$\lim_{\alpha\to\infty}{\boldsymbol{\beta}}^{\infty}_{\alpha,D}={\boldsymbol{\beta}}_{\ell_{2}}^{*}$ (128)
Although the same behaviour is observed as for the $D=2$ case, there are actually two main differences. The first one is that, for $D>2$, explicit regularization does not lead to a sparse bias. Indeed
@@
$R_{\alpha}({\boldsymbol{\beta}})\xrightarrow[]{\alpha\to 0}\lVert{\boldsymbol{\beta}}\rVert_{2/D},$ (130)